# DoFlow: Causal Generative Flows for Interventional and Counterfactual Time-Series Prediction

### v1

@@ 3 Interventional and Counterfactual Predictions
Figure 1: (A) RNN State Update. (B) Observational/Interventional Forecasting. Forecasts are generated by decoding from latent $z_{i,t}\sim N(0,1)$, conditioned on $\hat{H}_{i,t-1}$ updated with the last predicted $(\hat{x}_{i,t-1},\hat{x}_{\mathrm{pa}(i),t-1})$. (C) A factual observation $x_{i,t}^{\mathrm{F}}$ is encoded with its factual state $H_{i,t}^{\mathrm{F}}$ into $z_{i,t}^{\mathrm{F}}$, then decoded under the counterfactual state $\hat{H}_{i,t-1}^{\mathrm{CF}}$ to yield $\hat{x}_{i,t}^{\mathrm{CF}}$. Factual states $H_{i,t-1}^{\mathrm{F}}$ are updated from observed $x_{i,t-1}^{\mathrm{F}}$, while counterfactual states $\hat{H}_{i,t-1}^{\mathrm{CF}}$ are updated from the previously generated $\hat{x}_{i,t-1}^{\mathrm{CF}}$.
@@ 3.1 Observational and interventional prediction
2:Initialize hidden states $\hat{H}_{i,\tau}$ with $x_{i,1:\tau}$ for all $i=1,\dots,K$
@@
11:  $h_{i,t},h_{\mathrm{pa}(i),t}\xleftarrow[\text{}]{\text{update}}(\hat{x}_{i,t},\;\hat{x}_{\mathrm{pa}(i),t})$
@@ 3.2 Time series forecasting for counterfactual queries
Counterfactual forecasting follows the standard abduction–action–prediction procedure. Given a factual trajectory $\{x_{i,\tau+1:T}^{\mathrm{F}}\}_{i=1}^{K}$, we proceed as follows: (i) abduction – we infer latent variables by encoding each observed factual value into its latent representation $z_{i,t}^{\mathrm{F}}$ through the forward process (10), conditioned on a factual hidden state $H_{i,t-1}^{\mathrm{F}}$ that summarizes past factual observations; (ii) action – we apply the specified intervention schedule $\mathcal{I}$ with values $\{\gamma_{i,t}\}$ to the corresponding nodes; and (iii) prediction – we autoregressively generate the counterfactual trajectory $\{\hat{x}_{i,t}^{\mathrm{CF}}\}$ through the reverse process (11), starting from the abducted latent representations $\{z_{i,t}^{\mathrm{F}}\}$.
We first compute the factual hidden states $\{H^{\mathrm{F}}_{i,t}\}_{t=\tau}^{T-1}$ from the observed context $\{x_{i,1:\tau}\}$ and the observed factual trajectory $\{x^{\mathrm{F}}_{i,\tau+1:T}\}$. These factual states are used only for encoding factual values into their latent representations, which serve as the starting point for counterfactual decoding. At each time step $t$, nodes are predicted in topological order. For $(i,t)\in\mathcal{I}$, the counterfactual forecast is set as $\hat{x}_{i,t}^{\mathrm{CF}}:=\gamma_{i,t}$, and the counterfactual hidden state $\hat{H}_{i,t}^{\mathrm{CF}}$ is updated accordingly. For non-intervened nodes, the factual value $x_{i,t}^{\mathrm{F}}$ is first encoded as $z_{i,t}^{\mathrm{F}}=\Phi_{\theta}(x_{i,t}^{\mathrm{F}};\,H_{i,t-1}^{\mathrm{F}})$, then decoded under the counterfactual hidden state as $\hat{x}_{i,t}^{\mathrm{CF}}=\Phi_{\theta}^{-1}(z_{i,t}^{\mathrm{F}};\,\hat{H}_{i,t-1}^{\mathrm{CF}})$, followed by updating $\hat{H}_{i,t}^{\mathrm{CF}}$ with the generated value and its parents. The overall counterfactual generation procedure is summarized in Algorithm 2 and illustrated in Panel C of Figure 1.
@@
2:Obtain factual hidden states $\{H^{\mathrm{F}}_{i,t}\}_{t=\tau}^{T-1}$ from context $\{x_{i,1:\tau}\}$ and observed factual $\{x^{\mathrm{F}}_{i,\tau+1:T}\}$
3:Initialize counterfactual hidden states $\hat{H}^{\mathrm{CF}}_{i,\tau}=H_{i,\tau}$ with context $\{x_{i,1:\tau}\}$ for all $i=1,\dots,K$
@@
5:for $i=1,\dots,K$ do {nodes in topological order}
@@
12:  $h_{i,t},h_{\mathrm{pa}(i),t}\xleftarrow[\text{}]{\text{update}}(\hat{x}_{i,t}^{\mathrm{CF}},\;\hat{x}_{\mathrm{pa}(i),t}^{\mathrm{CF}})$
@@ 3.3 Additional property: likelihood-based anomaly detection
Another advantage of our framework is its ability to learn the density of the future trajectory over the forecasting window. For a single node (index $i$ omitted for clarity), we denote its density by $p_{\theta,X_{\tau+1:T}}(\cdot|\hat{H}_{\tau},z_{\tau+1:T})$, which represents the density of the future outcomes learned by the model, conditioned on the context state $\hat{H}_{\tau}$ and a sequence of simulated base samples $z_{\tau+1:T}$. The explicit form of the learned density, mapping from latent $z_{\tau+1:T}$ to data $\hat{x}_{\tau+1:T}$, is given as:
@@
$\small\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau},z_{\tau+1:T}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right].$ (12)
@@
Since anomalies deviate substantially from normal patterns, we expect that they have lower densities. Therefore, the log-density as computed in (12) naturally provides a principled measure of anomaly.
@@ 4 Theoretical Properties
In this section, we present a theoretical result on the counterfactual recovery of our algorithm under certain assumptions. For simplicity, we fix and omit the index $i$ from the notation. Recall that the encoding function is defined as $\Phi_{\theta}:\mathcal{X}\times\mathcal{H}\to\mathcal{Z}$ and the decoding function as $\Phi_{\theta}^{-1}:\mathcal{Z}\times\mathcal{H}\to\mathcal{X}$, corresponding to (10) and (11), respectively. Recall that the underlying structural causal model is given by $X_{t}:=f(X_{t-},X_{\text{pa},t-},U_{t})$. In the following, we present a supporting result on the counterfactual recovery properties of DoFlow. We begin by introducing the following assumptions.
@@
(A2) The structural causal equation $f(\cdot,U_{t})$ is monotone in $U_{t}$.
(A3) For the encoded latent variable $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$, the conditional distribution satisfies $p_{\theta}(Z_{t}\mid H_{t-1})=q(Z_{t})=N(Z_{t};0,1)$.
@@
Each node $X_{i}\in\mathbb{R}$, so (A2) is assumed in the univariate case. It is automatically satisfied under additive SCMs, i.e., $X_{t}=f^{*}(X_{t-},X_{\text{pa},t-})+U_{t}$. Under certain identifiability conditions, it also holds for non-linear models [70, 61]. Our theoretical results apply only to monotone SCM settings. One may notice that (A1)–(A2) mirror those in Bijective Generation Mechanisms (BGM) [47], which establish model-agnostic identifiability with an additional assumption on distribution matching. In contrast, our Corollary 4.5 provides model-specific, pointwise recovery for our CNF under (A1)–(A3) without requiring distribution matching; see Appendix D.
Note that (A3) implies that the encoded $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$ is statistically independent of $H_{t-1}$, and equivalently of $(X_{t-},X_{\text{pa},t-})$, in distribution. In the infinite-data limit with exact training, the continuous normalizing flow maps every $X_{t}$, conditioned on any fixed $(X_{t-},X_{\text{pa},t-})$, to the same base distribution $q(Z_{t})=N(0,1)$, so that (A3) holds exactly. In finite-sample regimes, deviations may occur; we empirically assess the validity of (A3) in Appendix E.1.
@@
Proposition 13 states that $Z_{t}$ is a function of the exogenous noise $U_{t}\sim\mathrm{Unif}[0,1]$. This assumption can be relaxed to other noise distributions. For example, if $Z\sim\mathcal{N}(0,1)$ with CDF $F$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any assignment $f(\cdot,U)$ can equivalently be written as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$. Therefore, the theorem’s conclusions also hold under Gaussian noise.
@@
$\small X^{\mathrm{CF}}_{t}\;=\;\begin{cases}\gamma_{t},&t\in\mathcal{I},\\[2.0pt] f\!\bigl(X^{\mathrm{CF}}_{t-},\,X^{\mathrm{CF}}_{\mathrm{pa},t-},\,U_{t}\bigr),&\text{otherwise,}\end{cases}$ (14)
where $U_{t}$ is abducted from the factual sample $X_{t}^{\mathrm{F}}$. In contrast, Algorithm 2 updates the estimated $\hat{H}_{t-1}^{\mathrm{CF}}$ recursively (Line 13), using the estimates $\hat{X}_{t-1}^{\mathrm{CF}}$ and $\hat{X}_{\mathrm{pa},t-1}^{\mathrm{CF}}$. We can now state the following result:
@@
Let Assumption 4.1 hold. Consider a factual sample generated by the structural causal model $X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and let its encoded latent be $Z_{t}^{\mathrm{F}}:=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}}\bigr)$. At time step $t$, we apply the intervention $\mathrm{do}\bigl(X_{t-}=\hat{X}_{t-}^{\mathrm{CF}},X_{\mathrm{pa},t-}=\hat{X}_{\mathrm{pa},t-}^{\mathrm{CF}}\bigr)$, yielding the counterfactual hidden state $\hat{H}_{t-1}^{\mathrm{CF}}$. Then the decoder recovers the true counterfactual at time step $t$ almost surely:
@@ 5 Experiments
In this section, we evaluate DoFlow on observational forecasting, causal forecasting under both interventional and counterfactual queries, and anomaly detection, using both synthetic and real-world datasets.
Because in real-world settings, the ground-truth counterfactual is never observable, and the ground-truth interventional is only observable if interventions are actively conducted under a correct causal DAG, we rely on synthetic experiments to obtain quantitative performance metrics. For real-world evaluation, we assess our model on interventional queries and anomaly detection using hydropower datasets from Argonne National Laboratory, and on interventional treatment effect estimation using the cancer-treatment dataset from [5].
@@ 5.1 Synthetic Data Experiments
We evaluate the model on multiple structurally diverse causal DAGs, using Root Mean Squared Error (RMSE) and Maximum Mean Discrepancy (MMD) as evaluation metrics. Since DoFlow generates a single counterfactual forecasting trajectory aimed at recovering the true counterfactual, MMD is not applicable. Therefore, we report only RMSE for counterfactual forecasting. Detailed simulation setups and metric definitions are provided in Appendix C.
For baseline comparisons in observational forecasting, we consider: a pure RNN-based method, the Gated Recurrent Unit (GRU) [10]; transformer-based methods, including the Temporal Fusion Transformer (TFT) [37] and the Time Series Dense Encoder (TiDE) [15]; and an all-MLP method with a specialized contextual mixing structure, the Time Series Mixer (TSMixer) [19]. We also compare against probabilistic methods, including DeepVAR [56] — a deep RNN-based model with multivariate Gaussian outputs, and MQF2 [29] — a convex deep neural network that learns multivariate quantile functions. We extend our acknowledgements to the Python packages Darts [24] and GluonTS [2], which we use to directly test several modern baselines in this paper.
To our knowledge, comparatively few works tackle forecasting interventional/counterfactual time-series on a causal DAG. Nevertheless, we can adapt strong observational forecasters for interventional simulations by training a separate model for each node in the causal DAG, conditioned on its parents. However, counterfactual generation remains challenging for these baselines.
Table 1: RMSE for observational, interventional, and counterfactual time series forecasting across causal structures: Tree, Diamond, and FC-Layer. Results for the Chain structure are provided in Table 6 (Appendix). MMD results are provided in Table 5 (Appendix). Reported values are averaged over 50 test batches, each containing 128 test series.
@@
DoFlow $\mathbf{0.57_{\pm.09}}$ $\mathbf{0.66_{\pm.11}}$ $\mathbf{0.65_{\pm.10}}$ $\mathbf{0.59_{\pm.13}}$ $\mathbf{0.68_{\pm.14}}$ $\mathbf{0.69_{\pm.14}}$ $0.52_{\pm.10}$ $\mathbf{0.71_{\pm.14}}$ $\mathbf{0.72_{\pm.14}}$ $\mathbf{0.31_{\pm.07}}$ $\mathbf{0.63_{\pm.14}}$ $\mathbf{0.69_{\pm.15}}$ $0.39_{\pm.12}$ $\mathbf{0.41_{\pm.13}}$ $\mathbf{0.49_{\pm.13}}$ $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.73_{\pm.15}}$ $\mathbf{0.64_{\pm.12}}$
@@
Tables 1 and 5 (Appendix) report the RMSE and MMD results for observational, interventional, and counterfactual time-series forecasting across multiple causal structures (detailed in Appendix C). DoFlow consistently delivers strong performance in standard observational forecasting and interventional forecasting compared with our adapted baselines, and uniquely supports counterfactual forecasting. We present one visual result of interventional and counterfactual in Figure 2.

### v2

@@ 3.1 Observational and interventional prediction
2:Initialize hidden states $\hat{H}_{i,\tau}$ with $(x_{i,1:\tau},x_{\mathrm{pa}(i),1:\tau})$ for all $i=1,\dots,K$
@@
11:  $h_{i,t}\leftarrow\text{RNN}(\hat{x}_{i,t},h_{i,t-1})$
@@
Figure 1: (A) RNN State Update. (B) Observational/Interventional Forecasting. Forecasts are generated by decoding from latent $z_{i,t}\sim N(0,1)$, conditioned on $\hat{H}_{i,t-1}$ updated with the last predicted $(\hat{x}_{i,t-1},\hat{x}_{\mathrm{pa}(i),t-1})$. (C) Counterfactual Forecasting. A factual observation $x_{i,t}^{\mathrm{F}}$ is encoded with its factual state $H_{i,t-1}^{\mathrm{F}}$ into $z_{i,t}^{\mathrm{F}}$, then decoded under the counterfactual state $\hat{H}_{i,t-1}^{\mathrm{CF}}$ to yield $\hat{x}_{i,t}^{\mathrm{CF}}$. Factual states $H_{i,t-1}^{\mathrm{F}}$ are updated from observed factual $(x_{i,t-1}^{\mathrm{F}},x_{\mathrm{pa}(i),t-1}^{\mathrm{F}})$, while counterfactual states $\hat{H}_{i,t-1}^{\mathrm{CF}}$ are updated from the previously predicted $(\hat{x}_{i,t-1}^{\mathrm{CF}},\hat{x}_{\mathrm{pa}(i),t-1}^{\mathrm{CF}})$.
@@ 3.2 Time series forecasting for counterfactual queries
Counterfactual forecasting follows the standard abduction–action–prediction procedure. Given a factual trajectory $\{x_{i,\tau+1:T}^{\mathrm{F}}\}_{i=1}^{K}$, we proceed as follows: (i) abduction – we infer latent variables by encoding each observed factual value into its latent representation $z_{i,t}^{\mathrm{F}}$ through the forward process (11), conditioned on a factual hidden state $H_{i,t-1}^{\mathrm{F}}$ that summarizes past factual observations; (ii) action – we apply the specified intervention schedule $\mathcal{I}$; and (iii) prediction – we generate the counterfactual trajectory $\hat{x}_{i,t}^{\mathrm{CF}}$ through the reverse process (12), starting from the abducted latent representations $z_{i,t}^{\mathrm{F}}$.
Specifically, we first compute the factual hidden states $\{H^{\mathrm{F}}_{i,t}\}_{t=\tau}^{T-1}$ for each $i$ from the observed context $\{x_{i,1:\tau}\}_{i=1}^{K}$ and the observed factual trajectory $\{x^{\mathrm{F}}_{i,\tau+1:T}\}_{i=1}^{K}$. These factual states are used only for encoding factual values into their latent representations, which serve as the starting point for counterfactual decoding. At each time-series step $t$, nodes are predicted in topological order. For $(i,t)\in\mathcal{I}$, the counterfactual forecast is set as $\hat{x}_{i,t}^{\mathrm{CF}}:=\gamma_{i,t}$, and its RNN state $h_{i,t}$ is updated accordingly. For non-intervened nodes, the factual value $x_{i,t}^{\mathrm{F}}$ is first encoded as $z_{i,t}^{\mathrm{F}}=\Phi_{\theta}(x_{i,t}^{\mathrm{F}};\,H_{i,t-1}^{\mathrm{F}})$, then decoded under the counterfactual hidden state as $\hat{x}_{i,t}^{\mathrm{CF}}=\Phi_{\theta}^{-1}(z_{i,t}^{\mathrm{F}};\,\hat{H}_{i,t-1}^{\mathrm{CF}})$, followed by updating $\hat{H}_{i,t}^{\mathrm{CF}}$. The overall counterfactual generation procedure is summarized in Algorithm 2 and illustrated in Panel C of Figure 1.
@@
2:Obtain factual hidden states $\{H^{\mathrm{F}}_{i,t}\}_{t=\tau}^{T-1}$ for each $i$ from context $\{x_{i,1:\tau}\}_{i=1}^{K}$ and observed factual $\{x^{\mathrm{F}}_{i,\tau+1:T}\}_{i=1}^{K}$
3:Initialize counterfactual hidden states $\hat{H}^{\mathrm{CF}}_{i,\tau}$ with context $(x_{i,1:\tau},x_{\mathrm{pa}(i),1:\tau})$ for all $i=1,\dots,K$
@@
5:for $i=1,\dots,K$ do {topological order}
@@
12:  $h_{i,t}\leftarrow\text{RNN}(\hat{x}_{i,t}^{\mathrm{CF}},h_{i,t-1})$
@@ 3.3 Additional property: likelihood-based anomaly detection
Another advantage of our framework is that it assigns an explicit log-density to its predicted future trajectory over the forecasting window. For a single node (index $i$ omitted for clarity), we denote this conditional density by $p_{\theta,X_{\tau+1:T}}(\cdot\mid\hat{H}_{\tau})$, where $\hat{H}_{\tau}$ is the context state. In DoFlow, $z_{\tau+1:T}$ are base samples used to generate the forecast $\hat{x}_{\tau+1:T}$. The explicit form of this learned density, mapping from latent $z_{\tau+1:T}$ to predicted $\hat{x}_{\tau+1:T}$, is given as:
@@
$\log p_{\theta,X_{\tau+1:T}}\!\left(\hat{x}_{\tau+1:T}\mid\hat{H}_{\tau}\right)=\sum_{t=\tau+1}^{T}\left[\log q\!\left(z_{t}\right)+\int_{0}^{1}\nabla\!\cdot\,v_{\theta}\bigl(x_{t}(s),s;\,\hat{H}_{t-1}\bigr)\,ds\right],$ (13)
where $\hat{x}_{t}=\Phi_{\theta}^{-1}(z_{t};\hat{H}_{t-1}),\ t\in\{\tau+1,\dots,T\}$.
@@
Since anomalies deviate substantially from normal patterns, we expect the model to assign lower density to its generated forecast trajectories when conditioned on anomalous contexts.
@@ 4 Theoretical Properties
Our analysis assumes that the underlying SCM on the given DAG satisfies the conditions in Assumption B.1 (see Appendix B.1), which are standard and fundamental in causal inference (Pearl, 2009; Hernán and Robins, 2010; Imbens and Rubin, 2015; Javaloy et al., 2023; Chao et al., 2024). These assumptions ensure that the interventional and counterfactual distributions we study are well-defined.
In this section, we present a theoretical result on the counterfactual recovery of our algorithm under certain assumptions. We fix a node and omit the index $i$ for simplicity. Recall that the encoding function is defined as $\Phi_{\theta}:\mathcal{X}\times\mathcal{H}\to\mathcal{Z}$ and the decoding function as $\Phi_{\theta}^{-1}:\mathcal{Z}\times\mathcal{H}\to\mathcal{X}$, corresponding to (11) and (12), respectively. Also, the underlying structural causal model is given by $X_{t}:=f(X_{t-},X_{\text{pa},t-},U_{t})$. In the following, we present a supporting result on the counterfactual recovery properties of DoFlow. We begin by introducing the following assumptions.
@@
(A2) The structural causal equation $f(\cdot,U_{t})$ is strictly monotone and continuous in $U_{t}$.
(A3) For the encoded latent variable $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$, the conditional distribution satisfies $p_{\theta}(Z_{t}\mid H_{t-1})=q(Z_{t})$.
@@
Each node $X_{t}\in\mathbb{R}$, so (A2) is assumed in the univariate case. It is automatically satisfied under additive SCMs, i.e., $X_{t}=f^{*}(X_{t-},X_{\text{pa},t-})+U_{t}$. Under certain identifiability conditions, it also holds for non-linear models (Zhang and Hyvarinen, 2012; Strobl and Lasko, 2023). Under (A2), for any fixed $(X_{t-},X_{\mathrm{pa},t-})$, the map $u\mapsto f(X_{t-},X_{\mathrm{pa},t-},u)$ is a bijection on the support of $U_{t}$. One may notice that (A1)–(A2) mirror those in Bijective Generation Mechanisms (BGM) (Nasr-Esfahany et al., 2023), which establish model-agnostic identifiability with an additional assumption on distribution matching. In contrast, our Corollary 4.5 provides model-specific, pointwise recovery for our CNF under (A1)–(A3) without requiring distribution matching; see Appendix B.2.
Note that (A3) implies that the encoded $Z_{t}=\Phi_{\theta}(X_{t};H_{t-1})$ is statistically independent of $H_{t-1}$, and equivalently of $(X_{t-},X_{\text{pa},t-})$, in distribution. In the infinite-data limit with exact training, the continuous normalizing flow maps every $X_{t}$, conditioned on any fixed $(X_{t-},X_{\text{pa},t-})$, to the same base distribution $q(Z_{t})=N(0,1)$, so that (A3) holds exactly.
@@
Proposition 14 states that $Z_{t}$ is a function of the exogenous noise $U_{t}$, and for convenience assumes $U_{t}\sim\mathrm{Unif}[0,1]$. This distributional assumption can be relaxed to other noise distributions. For example, if $Z\sim\mathcal{N}(0,1)$ with CDF $F$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any assignment $f(\cdot,U)$ can equivalently be written as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$.
@@
$X^{\mathrm{CF}}_{t}\;=\;\begin{cases}\gamma_{t},&t\in\mathcal{I},\\[2.0pt] f\!\bigl(X^{\mathrm{CF}}_{t-},\,X^{\mathrm{CF}}_{\mathrm{pa},t-},\,U_{t}\bigr),&\text{otherwise,}\end{cases}$ (15)
where $U_{t}$ is abducted from the factual sample $X_{t}^{\mathrm{F}}$. Also, $\hat{X}_{t-1}^{\mathrm{CF}}$ and $\hat{X}_{\mathrm{pa},t-1}^{\mathrm{CF}}$ denote the model predicted counterfactual values by Algorithm 2. We can now state the following result:
@@
Let Assumption 4.1 hold. Consider a factual sample generated by the structural causal model $X_{t}^{\mathrm{F}}=f(X_{t-},X_{\mathrm{pa},t-},U_{t})$, and let its encoded latent be $Z_{t}^{\mathrm{F}}:=\Phi_{\theta}\bigl(X_{t}^{\mathrm{F}};H_{t-1}^{\mathrm{F}}\bigr)$. At time step $t$, we obtain the counterfactual hidden state $\hat{H}_{t-1}^{\mathrm{CF}}$ from the predicted counterfactual past $(\hat{X}_{t-}^{\mathrm{CF}},\hat{X}_{\mathrm{pa},t-}^{\mathrm{CF}}\bigr)$. The decoder recovers the true counterfactual at time step $t$ almost surely:
$\hat{X}_{t}^{\mathrm{CF}}:=\Phi_{\theta}^{-1}\!\bigl(Z_{t}^{\mathrm{F}};\,\hat{H}_{t-1}^{\mathrm{CF}}\bigr)=X_{t}^{\mathrm{CF}}.$ (16)
@@ 5 Experiments
In this section, we evaluate DoFlow on observational forecasting, interventional and counterfactual forecasting, and anomaly detection, on both synthetic and real-world datasets. Our code is publicly available at https://github.com/StatFusion/DoFlow_Causal_Time_Series.
In real-world, the ground-truth counterfactual is never observable, and the ground-truth interventional is only observable if interventions are actively conducted under a correct causal DAG. Therefore, we rely on synthetic experiments to obtain quantitative performance metrics. For real-world evaluation, we assess our model on interventional queries and anomaly detection using hydropower datasets from Argonne National Laboratory, and on interventional treatment effect estimation using the cancer-treatment dataset from Bica et al. (2020a).
@@ 5.1 Synthetic Data Experiments
We evaluate the model across multiple structurally diverse causal DAGs using Root Mean Squared Error (RMSE), Maximum Mean Discrepancy (MMD), and Continuous Ranked Probability Score (CRPS). RMSE is reported for all settings. MMD and CRPS are computed only for observational and interventional forecasting, since counterfactuals in DoFlow yield a single deterministic trajectory. Detailed simulation setups and metric definitions are provided in Appendix D.
For baseline comparisons in observational forecasting, we consider: a pure RNN-based method, the Gated Recurrent Unit (GRU) (Che et al., 2018); transformer-based methods, including the Temporal Fusion Transformer (TFT) (Lim et al., 2021) and the Time Series Dense Encoder (TiDE) (Das et al., 2024); and an all-MLP method with a specialized contextual mixing structure, the Time Series Mixer (TSMixer) (Ekambaram et al., 2023). We also compare against probabilistic methods, including DeepVAR (Salinas et al., 2019) — a deep RNN-based model with multivariate Gaussian outputs, and MQF2 (Kan et al., 2022) — a convex deep neural network that learns multivariate quantile functions. We extend our acknowledgements to the Python packages Darts (Herzen et al., 2022) and GluonTS (Alexandrov et al., 2020), which we use to directly test several modern baselines in this paper.
To our knowledge, comparatively few works tackle interventional or counterfactual time-series forecasting on a causal DAG. Nevertheless, we can adapt strong observational forecasters for interventional task by training a separate model for each node in the causal DAG, conditioned on its parents; see Appendix E.1. However, counterfactual generation remains challenging for these baselines.
Table 1: RMSE for observational, interventional, and counterfactual time series forecasting across causal structures: Tree, Diamond, and FC-Layer. Results for the Chain structure are provided in Table 6 (Appendix). MMD and CRPS results are provided in Table 4 and Table 5 (Appendix), respectively. Reported values are averaged over 50 test batches, each containing 128 test series.
@@
DoFlow $\mathbf{0.57_{\pm.09}}$ $\mathbf{0.54_{\pm.10}}$ $\mathbf{0.11_{\pm.02}}$ $\mathbf{0.59_{\pm.13}}$ $\mathbf{0.62_{\pm.12}}$ $\mathbf{0.14_{\pm.02}}$ $0.55_{\pm.10}$ $\mathbf{0.57_{\pm.08}}$ $\mathbf{0.12_{\pm.01}}$ $\mathbf{0.31_{\pm.07}}$ $\mathbf{0.38_{\pm.08}}$ $\mathbf{0.19_{\pm.03}}$ $0.39_{\pm.12}$ $\mathbf{0.41_{\pm.13}}$ $\mathbf{0.16_{\pm.03}}$ $\mathbf{0.49_{\pm.08}}$ $\mathbf{0.36_{\pm.06}}$ $\mathbf{0.18_{\pm.04}}$
@@
Tables 1 (main) and 4, 5 (Appendix) report the RMSE, MMD, and CRPS results for observational, interventional, and counterfactual time-series forecasting across multiple causal structures (see Appendix D). DoFlow consistently delivers strong performance in standard observational forecasting and interventional forecasting compared with our adapted baselines, and uniquely supports counterfactual forecasting. We present visual results for both interventional and counterfactual forecasting in Figure 2 (main text), and in Figures 8 and 9 (Appendix).